the formual relating "n" (the number of cricket chirps per minute) to "t" ( the temperature in degrees fahrenheit) is "n"=4t-160.

A. Using the symbolic method, find how many times a cricket would chrip in a minute at 90 degrees fahrenheit.
B. It is evening, and a cricket is chirping 48 times per minute. Use a symbolic method to find the temperature.

Answer by checkley77(12844)

"n"=4t-160.
N=4*90-160
N=360-160
N=200 CRICKET CHIRPS PER MINUTE @ 90 DEGREES.
-------------------------------------------------
48=4T-160
4T=48+160
4T=208
T=208/4
T=52 IS THE TEMPERATURE FOR 48 CHIRPS PER MINUTE.